Greenhab Report – March 8th

Crew 206 Green Hab Report – 08-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (1 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 16.3°C

High: 30.4°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.3 gallons

Water in Blue Tank: 232.5 gallons

Crops watered at 0830 hours (2.15 gallons) and 1830 hours (2.15 gallons):


1) 08.30 AM 23.5° C / 37 %
2) 12.00 AM 26.9° C / 20 %
3) 15.00 PM 23.6° C / 20 %
4) 18.30 PM 20.4° C / 22 %

Changes to crops:
Here are the notable changes from the last week :
– One sunflower is in flower, an other one is also flowering
– The fenel planted on 15 JAN keeps growing (~ 3 cm)
– Still no sprouts from the tomatoes planted last week
– The cucumber is still very weak, almost dead
– The cress planted this week is now ~ 6 cm
– The sugar peas are still yellow, even with a bit of fertilizer
– Two lettuces are almost ready to harvest. The other are small (leafs
of ~ 2 cm)
– Microgreens and savory mix in tray almost ready to harvest

In the blue gardener :
– Basil, parsley, and sage are still abundant !
– Fenel is getting wider
– The tomatoe still has one sick branch
– Cilandro is now ~ 10 cm

Planted today :
– In a gardener : one row of Savory Mix, one row of Lettuce, and one row
of microgreens
– In a pot : chives

– Today I planted new crops in the greenhab : chives in a pot, and
savory mix + lettuces + microgreens in a gardener.

– I also made an inventory of the current plant in the greenhab and
their changes since last week.

– I opened the door 1 hour this morning, between 1000 and 1100.

– I put a bit of fertilizer on the tomatoe in the blue tray, and on
sugar peas.

Harvest: 8 g of basil and 1.6 g of sage

Support/supplies needed: None.

Journalist Report – March 8th

Crew 206 – 03/08/2019

Sol 12

Authors : Aurélien Mure, Astronomer and Executive Officier and Benjamin
Auzou, Journalist

"The Hab, our shelter"

After the incident of the weather station, we gave a second life and
role to the LOAC : the study of the air inside the Hab. Our goal is to
monitor the particles inside the Hab and compare with the data outside
the Hab. For its first hour of data collecting in the Hab, we submit the
LOAC to a massive particles number. As a matter of fact we cleaned the
Hab and so we raised a lot of dust from the ground. At the precise time
we wrote that report, the wind and the rain are threatening us; and we
are here, confined but above all protected in the Hab. We took advantage
of that day inside the Hab to precise our water monitoring. We spent a
lot of time evaluating the average volume of a flush in order to reduce
our unmeasured water volume.

As a crew fond of Escape Games, we played one before the mission as team
building, and it was a great success. Indeed, the crew has wonderful
skills to escape a room. Unfortunately, the Hab is not an Escape room
and we are here for 21 sols rather than an hour. Confinement is one of
the most interesting aspects of the simulation. It has many sneaky
impacts on every member of the crew. For example it is likely to have
some desires that we do not have in our everyday lives. We all really
wish to go for a hike with our parents, our friends although this is not
something we are used to do. Everybody is used to live far from his
relatives but when they miss us, we know that with a phone call, a train
ticket or a trip by car we can see them or talk to them in a couple of
hours. Here we are confined and isolated. It changes also our perception
of time. It is clear that the days are running really fast as we are in
our routine but we all have the strong impression that we are here for a
long time, not only two weeks. The moments in the day when we face
boredom feel much longer than they usually do. At our entry in the
station our motivation and will were extremely strong and even if we
enjoy the mission the little routine is slowly consuming our will to
begin new experiments or measures. The EVAs are a special moment. Even
if with the Simulation we can’t breathe the fresh air directly it is
satisfying to be able to walk out of the 8m of diameter Hab. As the
landscape is wonderful the EVAs are source of energy for every crew
member. As the wind is blowing around the station these days we are not
asking for EVAs as the Simulation is our first purpose. Indeed Martian
astronauts won’t go out if there is any risk outside so we decided to
stay in the Hab, maybe until Monday. Some of us can enjoy a small extra
space because of their role. In fact, Norbert can relax in the GreenHab,
Gaspard can work alone in the RAM and I (Aurélien) can observe the sun
in the Musk Observatory. Some of us also have activities to spend time :
books, mobile games, knitting, Rubik’s cube. We are spending our days
together, from 7:00 am to 10:00 pm so I want to remind the title of
yesterday’s Journalist Report, it is great to be an ‘unflappable crew’.

Operations Report – March 8th

Crew 206 Operations Report 08 – MAR – 2019

SOL: 12

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 12,3 hours

Solar— SOC % – Turned off at 79%, Turned on at 80% (18h30)

SOC tracking :
7h45: 78%
9h30: 74%
11h30: 98%
13h40: 92%
16h40: 86%
17h30: 82%
18h30: 80%

Diesel Reading – 65%
Propane Reading – 78%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 400 gallons

Water differential (static tank) – 9 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 232 gallons

Static to Loft Pump used – Yes

Water Meter: 0141699,3 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system tested.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Not Used

Hours: 78.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,9 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the
pipeline has a leak. Atila is working on it, more informations will follow.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used for Staff.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Still investigating. Nothing more to report today.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: None of the tarps seem to have suffered the
high winds today, everything else is doing fine too.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Sol Summary – March 8th

Crew 206 March 8th 2019
Sol 12

Summary Title :

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Since the begining of the week we are waking up
one or two minutes later every day, tomorrow we will have to get back in
track ! We still do our workout sessions as seriously as the first day
though, our scores have greatly increased. With the lack of an EVA we
decided to clean the Hab again, the dust settles everywhere so quickly
that it’s a necessity. We also had some long discussions on the methods
we use to measure the water consuption of flushes, we have to continue
averaging the water volume over many toilet flushes. For lunch we
prepared some pizzas, we ate a wonderfull lunch that made us all happy.
After that we played some Tarot, a popular french card game. In the
afternoon there was a TELEOP session, it worked fine, this experiment is
back on track ! As planned, there was also a VR session, we are
continuing to gather data.

Look ahead plan : We hope to go on EVA tomorrow but the weather doesn’t
seem promising, so we might just continue our indoor experiments in the
next days.

Anomalies in work : None

Weather : Cloudy all day with a few sunshines

Crew physical status : Fine

EVA : None

Reports to be filed : GreenHab, Journalist, Operation, EVA Request is
weather allows it

Support Requested : None

Copyright © The Mars Society. All rights reserved. | Main Site